How many liters of 7% acid must be mixed with how many liters of 12% acid solution by a chemist to obtian 6 liters of 10% acid?
Found 2 solutions by scott8148, Alan3354:
Answer by scott8148(6628)

(7%)(x) + (12%)(6 - x) = (10%)(6)

Answer by Alan3354(69443)

How many liters of 7% acid must be mixed with how many liters of 12% acid solution by a chemist to obtian 6 liters of 10% acid?
---------------
s = amount of 7%
t = amount of 12%
---------
s + t = 6 (total solution)
0.07s + 0.12t = 6*0.1 (total acid)
---
Solve for s & t